Group Selection Page
HELP
Controls Group
Data Analysis Group
ITDM
WP76
DET
Vacuum
EEE
WP78
WP74
FXE
SCS
SPB
MID
HED
SQS
SXP
Sample Environment
Operation
PSPO
XO
TS
Migrated2Zulip
|
Technical integration and coordination
SASE1 Status and SASE3 Readiness
|
Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther |
Not logged in |
 |
|
Message ID: 546
Entry time: 09 Sep 2024, 09:38
|
Author: |
Frederik Wolff-Fabris |
Location: |
General |
Severity: |
Information |
Subject: |
Notes XO Integration Meeting 09.09.2024 |
|
|
General Announcement:
- A. Galler informs a downtime of about 30min happened this morning in the North Branch due to a faulty vacuum controller.
- H. Sinn mentions on the beam drift in SA1: it was seen in both instruments over the weekend and is likely the comes from M1 but can be corrected by feedbacks in M2 and M3;
drift does not slowdown and together with XRO it will be looked if corrections on M1 can be applied tomorrow.
CTRL:
- Nothing to report.
DA:
- Nothing to report.
EEE:
- A motor in FXE attenuator had an issue on Saturday; it seems issue with encoders and to be further investigated.
VAC:
- Last week was quiet;
- Issue with ion pump controller behaving strangely; to be exchanged tomorrow. System is recovered and can run until tomorrow.
XPD:
- Nothing to report.
DET:
- Nothing major to report; no news on SPB AGIPD chiller.
ITDM:
- Nothing to report.
XRO:
- Nothing to report.
LAS:
- Nothing to report.
UNSYS:
- Issue with cell26/SA1 showing error in gap was fixed.
SPB:
- Nothing to report.
FXE:
- Beam drift was seen at FXE; it is communicated to PRC and XRO.
- Feedbacks for M2 and M3 manage to correct the beam drift.
MID:
- Nothing to report.
HED:
- Nothing to report.
SCS:
- Not present.
SQS:
- Nothing to report.
SXP:
- Not present. |